OpenStudy (anonymous):

Simplify the expression below. Squareroot of 6m over squareroot of 5 6m/5 squareroot of 6m squareroof of 30m over 5 squareroot of 30m over 25

OpenStudy (anonymous):

@Muzzack can you help me with this?

OpenStudy (anonymous):

@iGreen can you help me with this?

OpenStudy (anonymous):

and this one, square root of 4e over 3f

OpenStudy (anonymous):

\[\frac{ \sqrt{6m} }{ \sqrt{5} }?\]

OpenStudy (igreen):

D

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\sqrt{4e/3f}\]

OpenStudy (igreen):

Just graph them here: https://www.desmos.com/calculator The one that matches up with sqrt6m/sqrt5 is the answer.

OpenStudy (anonymous):

here are the answers

OpenStudy (anonymous):

\[=\frac{ \sqrt{6m} }{\sqrt{5} }*\frac{ \sqrt{5} }{ \sqrt{5} }=\frac{ \sqrt{30m} }{5 }\]

OpenStudy (anonymous):

okay I got that one right but this 2nd one i dont know.

OpenStudy (anonymous):

its the same rationalize the denominator

OpenStudy (anonymous):

@ghass1978 would it be B?

OpenStudy (anonymous):

yes

OpenStudy (igreen):

I'm 100% sure it's D.